# QUEST-A: Untrained Filtering with Trained Focusing led to Enhanced Quantum Architectures

### v1

@@ Abstract
Quantum architecture search (QAS) represents a fundamental challenge in quantum machine learning, with SOTA methods primarily categorized into training-free and gradient-guided approaches. However, treating QAS solely as either a discrete pruning process or a continuous optimization problem fails to balance accuracy and efficiency. This work decomposes QAS into two alternately solved subproblems: optimal circuit structure retrieval and parameter optimization. Based on this insight, we propose Quantum Untrained-Explored Synergistic Trained Architecture (QUEST-A), which implements rapid architecture pruning through inherent circuit properties and develops focused optimization with parameter reuse strategies. QUEST-A unifies discrete structure search and continuous parameter optimization within an evolutionary framework that integrates rapid pruning and fine-grained optimization. Experiments demonstrate QUEST-A’s superiority over existing methods: enhancing model expressivity in signal representation, maintaining high performance across varying complexities in image classification, and achieving order-of-magnitude precision improvements in variational quantum eigensolver tasks. These results validate QUEST-A’s effectiveness and provide transferable methodologies for QAS.
@@
Quantum machine learning , Quantum architecture search , Parameter reuse , Architecture pruning
@@ 1 Introduction
To address these challenges, Quantum architecture search (QAS) has emerged as a promising solution Wang et al. (2022); Du et al. (2022); Zhang et al. (2022); Wu et al. (2023); He et al. (2024b). QAS ingeniously combines classical machine learning optimization strategies with fundamental quantum physics principles, aiming to automatically discover optimal quantum circuit structures that satisfy learning task requirements under multiple constraints.
The core challenge of QAS lies in searching for optimal quantum circuit architectures that maintain expressive power while ensuring computational feasibility Lu et al. (2023); Ge et al. (2024). Currently, most QAS methods draw inspiration from the Neural Architecture Search (NAS) paradigm and can be broadly categorized into heuristic, differential-based, and sampling-based approaches based on their optimization strategies. Heuristic algorithms Wang et al. (2022); Williams and Gray (1998); Lamata et al. (2018); Yu et al. (2024) address the circuit architecture optimization problem through quadratic modeling and search the solution space according to predefined rules. These methods offer strong interpretability, providing clear insights into the optimization process. However, they suffer from limited scalability and can only probabilistically obtain global optima, often converging to local optima in large-scale problems. As an extension of variational quantum algorithms, differential-based QAS Zhang et al. (2022); Wu et al. (2023); Zhang et al. (2021); Du et al. (2022); Ma et al. (2024) constructs end-to-end differentiable workflows, utilizing gradient-guided searches with predefined supercircuits or modules. This approach effectively leverages gradient information; nevertheless, its search efficiency is constrained by preset structures and optimization techniques, resulting in computationally intensive training processes. Training-free QAS Anagolum et al. (2024); He et al. (2024b), an emerging efficient search mode derived from classical training-free architecture search methods Zhou et al. (2022), employs intrinsic characteristics of quantum circuits as performance proxy indicators to rapidly screen candidate circuit sets. The strength of this approach lies in its search efficiency, but rapid, undirected nature of the search also risks overlooking potentially superior architectural configurations.
@@
Considering the respective strengths and limitations of the aforementioned QAS methodologies, this research is motivated by the imperative to address the trade-off between expressibility and trainability in QAS. To this end, we propose the QUEST-A, which aims to surmount the inherent limitations of singular paradigms by synergistically combining the advantages of rapid untrained exploration and fine-grained trained optimization.
Specifically, QUEST-A adopts a multi-level search strategy, combining training-free rapid screening with variational quantum circuit fine-tuning optimization. It employs reupload circuits as a template for flexible architecture exploration. During the rapid screening phase, QUEST-A evaluates quantum circuits through complexity and expressivity indicators, which is coupled with subsequent gradient-based optimization to ensure search quality. The multi-level strategy naturally forms an iterative evolutionary process, where weight-sharing strategies accumulate performance advantages across generations, unifying discrete structure search and continuous parameter optimization. The contributions of our work are as follows:
@@
    Proposed the QUEST-A framework, integrating untrained rapid filtering with trained focusing strategies, overcoming the limitations of existing methods that overly emphasize either efficiency or accuracy, and achieving an optimized balance between search process efficiency and precision.
@@
    Viewed the natural evolution of the search process as an evolutionary sequence of candidate circuit ensembles, establishing rapid indexing and parameter transfer mechanisms. This approach achieved efficient transmission of advantageous parameters.
@@
    In numerical simulations and NISQ environment experiments across multiple types of tasks, QUEST-A demonstrated performance superior to state-of-the-art methods. Specifically, it achieved up to $81.8\%$ performance improvement in representation tasks, maintained over $75\%$ high accuracy in $10$-class image classification tasks, and enhanced precision by up to $11$ orders of magnitude in VQE tasks.
@@ 2.1 Quantum circuits for quantum machine learning
This process initiates with the encoding of classical data into quantum states, followed by the manipulation of these states through variational and entangling gates, and culminates in the conversion of results back to classical data via measurement. The training process of VQCs involves parameter optimization, typically employing gradient descent methods. Gradient computation in quantum circuits necessitates multiple circuit executions, with computational costs scaling linearly with the number of parameters. This characteristic, combined with the noise sensitivity of NISQ devicesPreskill (2018), underscores the significant impact of circuit structure on QML performance.
The influence of quantum circuit structure on learning efficacy is primarily manifested in three aspects: expressive power, training efficiency, and hardware compatibility. The intrinsic expressibility of a circuit is determined by the type, quantity, and arrangement of gates, independent of the parameter optimization process, and represents an inherent property of quantum circuits Sim et al. (2019). Training efficiency reflects the trade-off between expressibility and the number of trainable parameters -- complex circuits do not necessarily enhance expressibility but often increase learning costs. In the NISQ era, circuit design can also consider hardware connectivity constraints and noise characteristics.
@@ 2.2 Quantum Architecture Search
QAS is an optimization method aimed at finding optimal quantum circuit structures and parameters within a large-scale search space. The complexity of QAS can be quantified by analyzing its search space: as described in Fig. 2, consider a quantum circuit model comprising $n$ qubits and m layers, where each layer can select from a set of quantum gates $G$ with cardinality $k=|G|$. Under this model, the search space size reaches $O(k^{n×m})$ Ge et al. (2024). This exponential growth indicates that the computational complexity of QAS increases significantly as the number of qubits and circuit depth increase. Furthermore, QAS involves both discrete structure search and continuous parameter optimization Zhang et al. (2022), further augmenting the problem’s complexity.
@@
Based on this, the optimization objective of QAS can be formalized as $\min_{M\in G^{n\times m}}L(U,D),$ where $L$ denotes the loss function, and $D$ represents the dataset or target unitary matrix for the specific problem. The core of this optimization problem lies in identifying the quantum circuit structure M and its parameters that minimize the corresponding loss function for the specific problem within the $O(k^{n×m})$-scale search space. It is noteworthy that $G^{n\times m}$ represents the set of possible quantum gate combinations, the selection of which directly impacts the size and complexity of the search space.
@@ 2.3 Ansatz for QAS
The cascading CNOT circuit is currently the most commonly used ansatz McClean et al. (2018). By constructing complex entangled states, it establishes global correlations between qubits and serves as a prevalent template due to its universality. However, this structure assigns equal weights to all interactions, potentially obscuring feature importance Wang et al. (2022), and the complex entanglement patterns increase the difficulty of parameter optimization. The advancement of explainability of quantum computing Canatar et al. (2022); Cerezo et al. (2021); McClean et al. (2018) has given rise to various novel ansatzes. Hardware-efficient ansatzes Anagolum et al. (2024) consider the specific structure and gate sets of quantum hardware, while tensor network-inspired ansatzes Lazzarin et al. (2022) leverage theoretical insights from classical tensor networks.
Figure 3: Hybrid machine learning framework based on re-uploading quantum circuits. Data embedding becomes a repeatable and optimizable part of the variational circuit.
As an emerging quantum circuit ansatz, re-uploading circuits Pérez-Salinas et al. (2020) alternate data encoding and parameter variation processes, enabling multiple utilizations of input data and thus enhancing the circuit’s expressivity. This design establishes a complex, data-oriented non-unitary process that can approximate any continuous function Zhao et al. (2024), while providing increased flexibility and controllability in the search space. More importantly, by incorporating data encoding into the optimization scope, re-uploading circuits serve as an ideal template for QAS to potentially uncover superior problem representations Ono et al. (2023).
@@ 3.1 Overview of QUEST-A
Viewing QAS as a holistic optimization problem is challenging due to its large search space, suggesting the need for a decomposed optimization approach. A key insight of this study is that decomposing the QAS problem into two subproblems -- optimal circuit structure retrieval and optimal parameter optimization -- and solving these alternately is considerably simpler than addressing the overall optimization problem directly.
Based on this approach, QUEST-A optimizes quantum circuit structures through the synergistic effect of unsupervised exploration and supervised training. As shown in Fig. 4, the QUEST-A framework consists of three core components corresponding to untrained exploration, supervised training, and synergistic search strategies:
@@
[…]
@@
    Focused Architecture Optimization: QUEST-A applies variational quantum algorithms for fine-grained parameter optimization on pruned candidate structures. This process combines stochastic gradient descent for efficient parameter tuning. The introduced parameter reuse strategy transfers and utilizes learned parameter information across different structures through weight sharing.
@@
    Evolutionary Synthesis of Search Strategies: QUEST-A employs evolutionary algorithms as a meta-optimization framework, organically integrating rapid pruning and fine-tuned optimization processes. Population initialization corresponds to candidate generation in rapid architecture pruning, fitness evaluation combines untrained metrics and supervised training performance, selection operations are based on multi-objective optimization criteria, while crossover and mutation operations generate new candidate structures by manipulating quantum gate sequences and connection topologies.
These three coupled components form a closed-loop iterative optimization system. By alternating coarse-grained rapid pruning and fine-grained precise optimization within the evolutionary framework, QUEST-A achieves efficient, adaptive exploration of the quantum architecture space.
@@
3.2 Rapid Architecture Pruning
QAS can be considered an optimization problem, where appropriate reduction of the optimization space facilitates the optimization process. Rapid architecture pruning primarily focuses on discrete quantum circuit structures, thus further simplifying the problem into a discrete database retrieval task. Consequently, establishing an appropriate indexing mechanism is crucial to expedite the search process. Traditional supercircuit methods Wang et al. (2022); Du et al. (2022), typically construct and store the entire supercircuit at once as indexing , resulting in substantial memory overhead and training costs.
In contrast, QUEST-A adopts a dynamic indexing strategy, effectively overcoming these limitations. Formally, given the number of qubits $N$, maximum circuit depth $L$, and a set of available quantum gates $G(|G|=Q)$, the indexer $A$ for the search space $S$ can be defined as $A:{0,1,...,Q-1}^{N×L}\textrightarrow S$, where ${0,1,...,Q-1}^{N×L}$ represents the $N×L$-dimensional discrete index space. In the $t$-th iteration, the sampled specific quantum circuit structure $a^{t}$ can be expressed as $a^{t}=A(i^{t}),i^{t}\in{0,1,...,Q-1}^{N×L}$. This dynamic indexing method enhances memory efficiency by avoiding the need to store the entire supercircuit structure, while also improving computational efficiency by generating and evaluating specific structures only when necessary.
@@ 3.2 Rapid Architecture Pruning
The use of $P(c)$ aims to balance between circuit expressivity and trainability, selecting circuits that potentially avoid both under-parameterization and the barren plateau problem associated with over-parameterization, as suggested by studies Holmes et al. (2022); Cerezo et al. (2021) on quantum circuit complexity and trainability. This screening mechanism effectively eliminates candidate circuits that clearly do not meet task requirements. To complement the limitations of pure topological analysis, QUEST-A introduces an expressivity assessment $D(c)$ based on KL divergence. This method quantifies circuit expressivity by comparing the quantum state distribution generated by the circuit with the ideal Haar random distribution Jiang (2005). Specifically, for a quantum circuit c, its expressivity can be quantified by the following
$E(c)=-D_{KL}(P(c,F)||P_{Haar}(F))$ (2)
@@
The KL divergence metric captures subtle circuit expressivity characteristics that might be overlooked by simple topological analysis, providing potential performance estimates, especially when considering quantum properties such as entanglement and superposition. This approach aligns with work research Sim et al. (2019) on quantum circuit expressivity, emphasizing the importance of state distribution uniformity in evaluating circuit performance. In practical work, the relationship between KL divergence and candidate circuit performance does not exhibit a simple linear correlation. Instead, it provides a range of circuits with relatively stable performance, which can be used to filter out a series of circuit structures with lower training potential.
@@
3.3 Focused Architecture Optimization
After completing the rapid architecture pruning, the QUEST-A framework enters a more refined optimization phase. The primary objective of this stage is to efficiently optimize the parameters of the selected candidate quantum circuits. Through computationally intensive methods, this phase aims to fully exploit the potential of each circuit structure.

### v2

# Untrained Filtering with Trained Focusing for Superior Quantum Architecture Search

@@ Abstract
Quantum architecture search (QAS) represents a fundamental challenge in quantum machine learning. Unlike previous methods that treat it as a static search process, from a perspective on QAS as an item retrieval task in vast search space, we decompose the search process into dynamic alternating phases of coarse and fine-grained knowledge learning. We propose Quantum Untrained-Explored Synergistic Trained Architecture (QUEST-A),a framework through coarse-grained untrained filtering for rapid search space reduction and fine-grained trained focusing for precise space refinement in progressive QAS. QUEST-A develops an evolutionary mechanism with knowledge accumulation and reuse to enhance multi-level knowledge transfer in architecture searching. Experiments demonstrate QUEST-A’s superiority over existing methods: enhancing model expressivity in signal representation, maintaining high performance across varying complexities in image classification, and achieving order-of-magnitude precision improvements in variational quantum eigensolver tasks, providing a transferable methodology for QAS.
@@
Quantum machine learning , Quantum architecture search , Knowledge-driven search , Architecture pruning
@@ 1 Introduction
To address these challenges, Quantum architecture search (QAS) has emerged as a promising solution Wang et al. (2022); Du et al. (2022); Zhang et al. (2022); Wu et al. (2023); He et al. (2024b). Similar to Neural Architecture Search (NAS), QAS can be viewed as an item retrieval task within a vast architecture space, where the core challenge lies in efficiently retrieving target quantum circuit structures with limited prior knowledge of the search objective. This highlights the importance of dynamic knowledge accumulation during the search process.
Currently, most QAS methods draw inspiration from the NAS paradigm and can be broadly categorized into heuristic, differential-based, and sampling-based approaches based on their optimization strategies. Heuristic algorithms Wang et al. (2022); Williams and Gray (1998); Lamata et al. (2018); Yu et al. (2024) address the circuit architecture optimization problem through quadratic modeling and search the solution space according to predefined rules. These methods offer strong interpretability, providing clear insights into the optimization process. However, they suffer from limited scalability and can only probabilistically obtain global optima, often converging to local optima in large-scale problems. As an extension of variational quantum algorithms, differential-based QAS Zhang et al. (2022); Wu et al. (2023); Zhang et al. (2021); Du et al. (2022); Ma et al. (2024) constructs end-to-end differentiable workflows, utilizing gradient-guided searches with predefined supercircuits or modules. This approach effectively leverages gradient information; nevertheless, its search efficiency is constrained by preset structures and optimization techniques, resulting in computationally intensive training processes. Training-free QAS Anagolum et al. (2024); He et al. (2024b), an emerging efficient search mode derived from classical training-free architecture search methods Zhou et al. (2022), employs intrinsic characteristics of quantum circuits as performance proxy indicators to rapidly screen candidate circuit sets. The strength of this approach lies in its search efficiency, but rapid, undirected nature of the search also risks overlooking potentially superior architectural configurations. These methods essentially view QAS as a static search problem in a huge search space, using fixed retrieval strategies for either coarse-grained rapid pruning or fine-grained directed traversal, which explains their difficulty in balancing efficiency and accuracy.
@@
Considering the respective strengths and limitations of the aforementioned QAS methodologies, our work is motivated by the need to incorporate dynamic knowledge accumulation for progressive QAS. To this end, we propose the QUEST-A, which aims to surmount the limitations of singular paradigms by synergistically combining coarse-grained untrained filtering and fine-grained trained focusing. Specifically, QUEST-A iteratively combines coarse-grained untrained filtering for rapid search space reduction with fine-grained trained focusing for precise space refinement in a dynamic search process. Based on the flexibility of reupload circuits, the multi-level strategy naturally forms an iterative evolutionary process, where knowledge accumulation and reuse strategies unify coarse-grained rapid pruning and fine-grained precise search of the architecture space. This dynamic process emphasizes comprehensive knowledge utilization during search, avoiding the partial understanding of QAS in previous methods.
The contributions of our work are as follows:
@@
    Formulated QAS as an item retrieval task with dynamic knowledge learning, and developed a framework that iteratively combines coarse-grained untrained filtering for rapid search space reduction with fine-grained trained focusing for precise space refinement.
@@
    Developed an evolutionary mechanism with knowledge accumulation and reuse strategies, integrating coarse-grained untrained filtering and fine-grained trained focusing through genetic operations and fitness-guided selection, enabling knowledge transfer and accumulation during searching.
@@
    In numerical simulations and NISQ environment experiments across multiple types of tasks, QUEST-A demonstrated performance superior to state-of-the-art methods. Specifically, it achieved up to $81.8\%$ performance improvement in representation tasks, maintained over $85\%$ high accuracy in $10$-class image classification tasks, and enhanced precision by up to $11$ orders of magnitude in VQE tasks.
@@ 2.1 Quantum Circuits for Quantum Machine Learning
This process initiates with the encoding of classical data into quantum states, followed by the manipulation of these states through variational and entangling gates, and culminates in the conversion of results back to classical data via measurement. The training process of VQCs involves parameter optimization, typically employing gradient descent methods. From a knowledge learning perspective, gradient computation in quantum circuits provides fine-grained knowledge about circuit parameters, costs scaling linearly with the number of parameters. This characteristic, combined with the noise sensitivity of NISQ devicesPreskill (2018), underscores the significant impact of circuit structure on QML performance.
The influence of quantum circuit structure on learning performance is primarily manifested in three aspects: circuit scale, training potential, and hardware adaptability. The circuit scale, one of the most concerned properties in the NISQ era, is determined by the type, quantity, and arrangement of gates, which can be evaluated through inherent characteristics Sim et al. (2019). Training potential reflects the trade-off between expressibility and the number of trainable parameters -- complex circuits do not necessarily enhance expressibility but often increase learning costs. In the NISQ era, circuit design can also consider hardware connectivity constraints and noise characteristics.
@@ 2.2 Quantum Architecture Search
QAS is an item retrieval task aimed at finding optimal quantum circuit structures and parameters within a large-scale search space. The complexity of QAS can be quantified by analyzing its search space: as described in Fig. 2, consider a quantum circuit model comprising $n$ qubits and m layers, where each layer can select from a set of quantum gates $G$ with cardinality $k=|G|$. Under this model, the search space size reaches $O(k^{n×m})$ Ge et al. (2024). This exponential growth indicates that the computational complexity of QAS increases significantly as the number of qubits and circuit depth increase. Furthermore, QAS involves both discrete structure search and continuous parameter optimization Zhang et al. (2022), further augmenting the problem’s complexity.
@@
based on this, the search objective of QAS can be formalized as $\min_{M\in G^{n\times m}}L(U,D),$ where $L$ denotes the loss function, and $D$ represents the dataset or target unitary matrix for the specific problem. The core of QAS lies in efficiently retrieving the quantum circuit structure $M$ and its parameters that minimize the corresponding loss function within the $O(k^{n×m})$-scale search space with limited prior knowledge. It is noteworthy that $G^{n\times m}$ represents the set of possible quantum gate combinations, the selection of which directly impacts the size and complexity of the search space.
@@ 2.3 Ansatz for QAS
The cascading CNOT circuit is currently the most commonly used ansatz McClean et al. (2018). By constructing complex entangled states, it establishes global correlations between qubits and serves as a prevalent template due to its universality. However, this structure assigns equal weights to all interactions, potentially obscuring feature importance Wang et al. (2022), and the complex entanglement patterns increase the difficulty of parameter optimization. The advancement in understanding quantum circuit characteristics Canatar et al. (2022); Cerezo et al. (2021); McClean et al. (2018) has given rise to various novel ansatzes. Hardware-efficient ansatzes Anagolum et al. (2024) consider the specific structure and gate sets of quantum hardware, while tensor network-inspired ansatzes Lazzarin et al. (2022) leverage theoretical insights from classical tensor networks.
Figure 3: Hybrid machine learning framework based on re-uploading quantum circuits.
As an emerging quantum circuit ansatz, re-uploading circuits Pérez-Salinas et al. (2020) alternate data encoding and parameter variation processes, enabling multiple utilizations of input data and thus enhancing the circuit’s expressivity. This design establishes a complex, data-oriented non-unitary process that can approximate any continuous function Zhao et al. (2024), while providing increased flexibility and controllability in the search space. More importantly, by incorporating data encoding into the optimization scope Ono et al. (2023), re-uploading circuits’ modular and repeatable structure provides an ideal template for efficient knowledge transfer and accumulation during architecture searching.
@@ 3.1 Overview of QUEST-A
Traditional QAS approaches rely on fixed search strategies and face inherent difficulties in balancing search accuracy and efficiency. This limitation suggests the need for a dynamic knowledge-driven approach. A key insight of this study is the decomposition of QAS into two alternating phases: coarse-grained retrieval guided by circuit inherent properties, and fine-grained retrieval guided by task performance. This decomposition strategy provides a more tractable solution compared to directly addressing the entire search problem. As shown in Fig. 4, the QUEST-A framework consists of three core components corresponding to untrained filtering, trained focusing, and knowledge transfer strategies:
@@
    Untrained Filtering: This component constructs the initial search space using re--uploading circuit templates to indexes and manages candidate structures through supercircuit technology. The pruning process employs two complementary coarse--grained evaluation methods: first analyzing structural complexity using Directed Acyclic Graph (DAG) topological features, then calculating the expressive power of quantum state distributions via KL-divergence.
@@
    Trained Focusing: QUEST-A applies fine-grained search on filtered candidate spaces through gradient-guided mechanisms. This process progressively refines promising architectures in the reduced search space. The knowledge reuse strategy enables transfer of architectural knowledge between candidates, guiding the search towards superior regions.
@@
    Evolutionary Synthesis of Knowledge Transfer: QUEST-A employs evolutionary algorithms as a knowledge transfer framework, where each generation represents a specific region of the architecture space. Population initialization corresponds to candidate search regions generation in untrained filtering, fitness evaluation combines coarse-grained circuit properties and fine-grained task performance, selection operations are based on multi-level knowledge criteria, while crossover and mutation operations generate new search regions through structure manipulation.
These three coupled components form a closed-loop search system. By alternating coarse-grained untrained filtering and fine-grained trained focusing within the evolutionary framework, QUEST-A achieves architecture discovery with progressive knowledge accumulation.
@@
3.2 Untrained Filtering
QAS can be considered a retrieval task, where appropriate reduction of the search space facilitates the knowledge learning process. Untrained filtering primarily focuses on discrete quantum circuit structures, thus further simplifying the problem into a discrete database retrieval task. Consequently, establishing an appropriate indexing mechanism is crucial to enable efficient knowledge discovery. Traditional supercircuit methods Wang et al. (2022); Du et al. (2022), typically construct and store the entire supercircuit at once as indexing , resulting in substantial memory overhead and training costs.
In contrast, QUEST-A adopts a dynamic indexing strategy, effectively overcoming these limitations. Formally, given the number of qubits $N$, maximum circuit depth $L$, and a set of available quantum gates $G(|G|=Q)$, the indexer $A$ for the search space $S$ can be defined as $A:\{0,1,...,Q-1\}^{N×L}\rightarrow S$, where $\{0,1,...,Q-1\}^{N×L}$ represents the $N×L$-dimensional discrete index space. In the $t$-th iteration, the sampled specific quantum circuit structure $a^{t}$ can be expressed as $a^{t}=A(i^{t}),i^{t}\in\{0,1,...,Q-1\}^{N×L}$. This dynamic indexing method enhances memory efficiency by avoiding the need to store the entire supercircuit structure, while also improving computational efficiency by generating and evaluating specific structures only when necessary.
@@ 3.2 Untrained Filtering
The use of $P(c)$ aims to balance between circuit expressivity and trainability, selecting circuits that potentially avoid both under-parameterization and the barren plateau problem associated with over-parameterization, as suggested by studies Holmes et al. (2022); Cerezo et al. (2021) on quantum circuit complexity and trainability. This screening mechanism effectively eliminates candidate circuits that clearly do not meet task requirements. To complement the coarse-grained topological analysis, QUEST-A introduces an expressivity assessment $D(c)$ based on KL divergence. This method quantifies circuit expressivity by comparing the quantum state distribution generated by the circuit with the ideal Haar random distribution Jiang (2005). Specifically, for a quantum circuit c, its expressivity can be quantified by the following:
$E(c)=-D_{KL}(P(c,F)||P_{Haar}(F)),$ (2)
@@
Topological analysis, though effective for circuit complexity evaluation, cannot capture quantum-specific properties of circuits. KL divergence complements this limitation by providing a quantitative measure of circuit expressivity. This approach aligns with work research Sim et al. (2019) on quantum circuit expressivity, emphasizing the importance of state distribution uniformity in evaluating circuit performance. In practical work, the relationship between KL divergence and candidate circuit performance does not exhibit a simple linear correlation. Instead, it provides a range of circuits with relatively stable performance, which can be used to filter out a series of circuit structures with lower training potential.
@@
3.3 Trained Focusing
After completing the untrained filtering, the QUEST-A framework enters the trained focusing phase. The primary objective of this stage is to perform fine-grained search in the reduced architecture space. Through computationally intensive methods, this phase aims to fully exploit the potential of each circuit structure.